# CPHT-RR081.112024, December 2024

### v1

@@ 2 Classical minisuperspace model
Let us consider four-dimensional cosmologies, in the presence of a positive cosmological constant $\Lambda$ and in the absence of matter. In General Relativity, this can be described from the Lorentzian action
@@
For a closed, homogeneous and isotropic Universe, space is a 3-sphere and the metric takes the following Friedmann-Lemaître-Robertson-Walker form
@@
where ${\cal A}$ is an arbitrary positive, invertible function of a new field $q$. In terms of $q(x^{0})$, the classical Lagrangian in Eq. (2.3) becomes
@@
is the effective potential. [^2] Note that $\gamma_{{\rm q}{\rm q}}$ is negative. The kinetic term is therefore of the opposite sign to that of a scalar field. In general, the target space of a minisuperspace model comprising more than one degree of freedom is Lorentzian [1]. The conjugate momentum of the field $q$ is given by
@@ 3 Ambiguities of the canonical-quantization
This algebra can be represented by considering a Hilbert space of wavefunctions that describe the possible states of the Universe. Since these functions depend on the coordinate ${\rm q}$ of the target space ${\cal T}$, the commutation relation is achieved as usual by substituting
@@
Using Eqs. (2.13), (3.14) and (3.16), this contraint becomes the WDW equation for $\psi({\rm q})$.
@@ 4 Ambiguity of the path-integral quantization
$\begin{split}\psi({\rm q})&=\int_{0}^{+\infty}\!\!{\rm d}\ell\int_{{\rm q}_{\rm i}}^{{\rm q}}{\cal D}q\;e^{{i\over\hbar}\int_{0}^{1}{\rm d}x^{0}\,L(\ell,q,\dot{q})}\,,\\ &=\int_{0}^{+\infty}{\rm d}\ell\int_{\check{\rm q}_{\rm i}}^{\check{\rm q}}{\cal D}\check{q}\;{\cal J}\,e^{{i\over\hbar}\int_{0}^{1}{\rm d}x^{0}\,\check{L}(\ell,\check{q},\dot{\check{q}})}\,,\end{split}$ (4.33)
@@ 5 Inner product of the Hilbert space and probability amplitudes
Indeed, using only the definition (5.47)—i.e. without knowing what the expression of $\mu$ is—it is straightforward to show that the WDW equation (3.22) can be written as
@@
$\begin{split}Y^{\rm q}&=\gamma^{{\rm q}{\rm q}}\,{\rho^{{\rm q}\prime}\over\rho^{\rm q}}-{(\gamma^{{\rm q}{\rm q}}\sqrt{-\gamma\mu})^{\prime}\over\sqrt{-\gamma\mu}}-\gamma^{{\rm q}{\rm q}}\,{\nabla_{\!{\rm q}}\sqrt{\mu}\over\sqrt{\mu}}\,,\\ W&=\omega-{\nabla^{2}\!\sqrt{\mu}\over\sqrt{\mu}}-Y^{\rm q}\,{\nabla_{\!{\rm q}}\sqrt{\mu}\over\sqrt{\mu}}\,.\end{split}$ (5.50)
Thanks to the particular form of $\mu$ given in Eq. (5.45), the expression of $Y^{\rm q}$ reduces to
$Y^{\rm q}=i\nabla^{\rm q}\theta\,,$ (5.51)
which is pure imaginary and explicitly a vector. Using the expression of $\operatorname{Im}\omega$ given in Eq. (5.46), the extra potential term of the equation becomes
@@ 6 Determination of the ambiguities $\rho^{\rm q}$, $\omega$ and uniqueness of the quantum theory
${\cal A}_{0}(q_{0})=\left({3\over 2}\,{1\over\sqrt{6v_{3}}}\right)^{3\over 2}|q_{0}-{\rm q}_{0*}|^{2\over 3}\,,$ (6.60)
@@
$\begin{split}&i\hbar\,{U_{0}^{(n-1)}({\rm q}_{0{\rm i}},{\rm q}_{0},\ell-\varepsilon)-U_{0}^{(n)}({\rm q}_{0{\rm i}},{\rm q}_{0},\ell)\over-\varepsilon}=\\ &\,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\hbar^{2}\over 2}\,{1\over J({\rm q}_{0})}\,{\partial^{2}\over\partial q_{0(n-1)}^{2}}\big{[}J(q_{0(n-1)})\,U_{0}^{(n-1)}({\rm q}_{0{\rm i}},q_{0(n-1)},\ell-\varepsilon)\big{]}\Big{|}_{q_{0(n-1)={\rm q}_{0}}}\\ &\,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +V_{0}({\rm q}_{0})\,U_{0}^{(n-1)}({\rm q}_{0{\rm i}},{\rm q}_{0},\ell-\varepsilon)\,.\end{split}$ (6.77)
@@
$q\longrightarrow\widehat{\rm q}\,,\qquad-i\hbar\,\frac{\partial}{\partial{\rm q}}\longrightarrow\widehat{\pi}_{q}$ (6.85)
@@ 7 Comparison with previous works
where ${\cal R}$ is the Ricci scalar of the $m$-dimensional target space ${\cal T}$, and $\xi$ is a constant. Notice however that this form is not the most general one involving only terms with at most two derivatives. Among many possibilities, a term $\xi^{\prime}{\cal R}V\psi$ for any constant $\xi^{\prime}$, or a term involving any function of the potential, i.e. $f(V)\psi$ rather than $V\psi$, are allowed. Moreover, Ref. [10] imposes that Eq. (7.116) and its solutions are proportional to their counterparts, when one changes the lapse function $N$ to $\tilde{N}=\gamma^{2}\tilde{N}$, where $\gamma$ is an arbitrary function of the $m$ degrees of freedom. The idea is that the resulting equation and wavefunctions should give a physical description equivalent to the initial one. It is then claimed that this constraint determines what the value of $\xi$ is. However, nowhere in the derivation, the bounds of the integrals which define the actions are specified and the issues raised in the two previous paragraphs are ubiquitous. Moreover, as explained in our work, to conclude that a priori different Hilbert spaces of wavefunctions are equivalent, their inner products must be equal. However, inner products are not discussed in Ref. [10].

### v2

@@ 2 Classical minisuperspace model
Let us consider four-dimensional cosmologies, in the presence of a positive cosmological constant $\Lambda$ and in the absence of matter. [^1] The four-dimensional pure Einstein theory with vanishing cosmological constant does not yield non-trivial homogeneous and isotropic cosmological solutions with compact closed spatial sections and constant positive curvature. Non-isotropic homogeneous solutions with flat toroidal spatial sections do exist, but to realize these we need to retain more than one metric degree of freedom in the minisuperspace analysis. In this work, we will analyze only one-dimensional minisuperspace models. In General Relativity, this can be described from the Lorentzian action
@@
For a closed, homogeneous and isotropic Universe, space is a 3-sphere and the metric takes the following Friedmann-Lemaître-Robertson-Walker form [^3] It is essential for our formalism to consider spatially compact cosmological models, for which the minisuperspace Lagrangian and action are finite, and the formulation of the wavefunction of the universe in terms of a path integral is possible. If the spatial sections were non-compact, the minisuperspace action would be infinite, as this is proportional to the volume of the spatial sections, and the wavefunction path integral could not be defined. Flat or open homogeneous and isotropic universes with non-compact spatial sections of vanishing and constant negative curvature, respectively, cannot be considered. The open cases (for which the spatial sections have constant negative curvature) would require non-trivial compactifications and boundary conditions, breaking isotropy and possibly homogeneity. So we do not consider these in this work that focusses on simple one-dimensional minisuperspace models. Also, when the cosmological constant is vanishing, non-trivial homogeneous flat four-dimensional cosmological solutions, with compact toroidal sections, cannot be realized in the context of simple one-dimensional minisuperspace models. For positive cosmological constant, we could consider the case of a homogeneous model with a single minisuperspace degree of freedom (single scale factor degree of freedom) and flat toroidal spatial sections, breaking global isotropy. In this case, all of our conclusions and methods concerning the ordering ambiguities presented in the next section apply since, as we will argue, the kinetic term in the Hamiltonian remains unaffected.
@@
Note that this one-dimensional minisuperspace model does not contain any propagating physical degrees of freedom, and the full solution at the classical level can be obtained by solving a first-order constraint equation (obtained after fixing time-reparametrization invariance). The equation of motion for the scale factor follows by differentiating this constraint equation. See Ref. [23] for further discussions.
@@
where ${\cal A}$ is an arbitrary positive, invertible function of a new field $q$. Such field redefinitions of the scale factor may prove useful and practical in finding solutions of the WDW equation at the quantum level. In addition, they are linked with ambiguities in the path-integral quantization of the model, as we explain explicitly in Sect. 4. In terms of $q(x^{0})$, the classical Lagrangian in Eq. (2.3) becomes
@@
is the effective potential. [^4] Note that $\gamma_{{\rm q}{\rm q}}$ is negative. The kinetic term is therefore of the opposite sign to that of a scalar field. In general, the target space of a minisuperspace model comprising more than one degree of freedom is Lorentzian [1]. It is well known that due to gauge symmetries, pure Einstein gravity does not yield any propagating ghost. The field $q$ is a background mode that does not correspond to a propagating degree of freedom. However, due to time-reparametrization invariance, the Hamiltonian is constrained to vanish and the only allowable solutions are the ones corresponding to zero energy. No large negative energy solution is allowed. The conjugate momentum of the field $q$ is given by
@@ 3 Ambiguities of the canonical-quantization
This algebra can be represented by considering a Hilbert space of wavefunctions that describe the possible states of the Universe. Since these functions depend on the coordinate ${\rm q}$ of the target space ${\cal T}$, the commutation relation is achieved as usual by substituting [^5] Notice that the inverse minisuperspace metric appearing in the kinetic term of the Hamiltonian may yield powers of the inverse of the position operator $\hat{q}$. As usual, the domain of the position operator $\hat{q}$ can be considered to be the larger space of tempered distributions and its eigenfunctions the Dirac $\delta$-function distributions centered at the coordinate positions ${\rm q}$ (with eigenvalues ${\rm q}$). These eigenfunctions are non-normalizable. Notice that since $a={\cal{A}}({\rm q})\geq 0$, we have $(\operatorname{sign}{{\cal A}})({\rm q}-{\rm q}_{*})\geq 0$, where ${\rm q}_{*}={\cal A}^{-1}(0)$ and $\operatorname{sign}{{\cal A}}$ is constant since ${\cal A}$ is invertible. The constraint $(\operatorname{sign}{{\cal A}})({\rm q}-{\rm q}_{*})\geq 0$ can be implemented by using a projection operator $\hat{\Pi}=\Theta((\operatorname{sign}{{\cal A}})({\rm q}-{\rm q}_{*}))$, where $\Theta$ is the Heaviside step function. If $(\operatorname{sign}{\cal A})\,{\rm q}_{*}>0$, the position operator is invertible along the half line $(\operatorname{sign}{{\cal A}})({\rm q}-{\rm q}_{*})\geq 0$. Otherwise, we have $(\operatorname{sign}{\cal A})\,{\rm q}_{*}\leq 0$. In this second case, the inverse position operator is singular at ${\rm q}=0$ and is thus defined for $(\operatorname{sign}{{\cal A}})({\rm q}-{\rm q}_{*})\geq 0$, ${\rm q}\neq 0$. However, despite the singularity at ${\rm q}=0$, the wavefunction solution of the WDW equation and the corresponding probability density can be well behaving there, resolving this singularity. In Ref. [18], we show that this is indeed the case.
@@
Using Eqs. (2.13), (3.14) and (3.16), this constraint becomes the WDW equation for $\psi({\rm q})$.
@@ 4 Ambiguity of the path-integral quantization
$\begin{split}\psi({\rm q})&=\int_{0}^{+\infty}\!\!{\rm d}\ell\int_{{\rm q}_{\rm i}}^{{\rm q}}{\cal D}q\;e^{{i\over\hbar}\int_{0}^{1}{\rm d}x^{0}\,L(\ell,q,\dot{q})}\,,\\ &=\int_{0}^{+\infty}{\rm d}\ell\int_{\check{\rm q}_{\rm i}}^{\check{\rm q}}{\cal D}\check{q}\;{\cal J}\,e^{{i\over\hbar}\int_{0}^{1}{\rm d}x^{0}\,\check{L}(\ell,\check{q},\dot{\check{q}})}=\psi({\rm q}(\check{\rm q}))\,,\end{split}$ (4.33)
@@ 5 Inner product of the Hilbert space and probability amplitudes
Indeed, using only the definition (5.47)—i.e. without knowing what the expression of the scalar $\mu$ is—it is straightforward to show that the WDW equation (3.22) can be written as
@@
$\begin{split}Y^{\rm q}&=\gamma^{{\rm q}{\rm q}}\,{\left(\displaystyle{\rho^{\rm q}\over-\gamma^{{\rm q}{\rm q}}\sqrt{-\gamma}\mu}\right)^{\prime}\over\displaystyle{\rho^{\rm q}\over-\gamma^{{\rm q}{\rm q}}\sqrt{-\gamma}\mu}}={\nabla^{\rm q}\!\left(\displaystyle{\rho^{\rm q}\over-\gamma^{{\rm q}{\rm q}}\sqrt{-\gamma}\mu}\right)\over\displaystyle{\rho^{\rm q}\over-\gamma^{{\rm q}{\rm q}}\sqrt{-\gamma}\mu}}\end{split}$ (5.50)
and
$W=\omega-{\nabla^{2}\!\sqrt{\mu}\over\sqrt{\mu}}-Y^{\rm q}\,{\nabla_{\!{\rm q}}\sqrt{\mu}\over\sqrt{\mu}}\,.$ (5.51)
In the second expression for $Y^{\rm q}$, we have replaced the derivative by a covariant derivative, which is justified by the fact that the quantity in parenthesis is a contravariant vector density of weight $-1$. As a result, $Y^{\rm q}$ is a vector and $W$ a scalar under reparametrizations of the target space ${\cal T}$. Thanks to the particular form of $\mu$ given in Eq. (5.45), the expression of $Y^{\rm q}$ reduces to the purely imaginary quantity
$Y^{\rm q}=i\nabla^{\rm q}\theta\,.$ (5.52)
Using the expression of $\operatorname{Im}\omega$ given in Eq. (5.46), the extra potential term of the equation becomes
@@ 6 Determination of the ambiguities $\rho^{\rm q}$, $\omega$ and uniqueness of the quantum theory
${\cal A}_{0}(q_{0})=\left({3\over 2}\,{1\over\sqrt{6v_{3}}}\right)^{2\over 3}|q_{0}-{\rm q}_{0*}|^{2\over 3}\,,$ (6.61)
@@
$\begin{split}&i\hbar\,{U_{0}^{(n-1)}({\rm q}_{0{\rm i}},{\rm q}_{0},\ell-\varepsilon)-U_{0}^{(n)}({\rm q}_{0{\rm i}},{\rm q}_{0},\ell)\over-\varepsilon}=\\ &\,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\hbar^{2}\over 2}\,{1\over J({\rm q}_{0})}\,{\partial^{2}\over\partial q_{0(n-1)}^{2}}\big{[}J(q_{0(n-1)})\,U_{0}^{(n-1)}({\rm q}_{0{\rm i}},q_{0(n-1)},\ell-\varepsilon)\big{]}\Big{|}_{q_{0(n-1)={\rm q}_{0}}}\\ &\,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +V_{0}({\rm q}_{0})\,U_{0}^{(n-1)}({\rm q}_{0{\rm i}},{\rm q}_{0},\ell-\varepsilon)+\cal O(\varepsilon)\,.\end{split}$ (6.78)
@@
${\rm q}\longrightarrow\widehat{q}\,,\qquad-i\hbar\,\frac{\partial}{\partial{\rm q}}\longrightarrow\widehat{\pi}_{q}$ (6.86)
@@ 7 Comparison with previous works
where ${\cal R}$ is the Ricci scalar of the $m$-dimensional target space ${\cal T}$, and $\xi$ is a constant. Notice however that this form is not the most general one involving only terms with at most two derivatives. Among many possibilities, a term $\xi^{\prime}{\cal R}V\psi$ for any constant $\xi^{\prime}$, or a term involving any function of the potential, i.e. $f(V)\psi$ rather than $V\psi$, are allowed. Moreover, Ref. [10] imposes that Eq. (7.117) and its solutions are proportional to their counterparts, when one changes the lapse function $N$ to $\tilde{N}=\gamma^{2}\tilde{N}$, where $\gamma$ is an arbitrary function of the $m$ degrees of freedom. [^17] For a more complete discussion concerning this symmetry and invariance under conformal rescalings of the minisuperspace metric in the higher dimensional cases, see Ref. [24]. Such a symmetry may play a role in fixing ordering ambiguities in the higher-dimensional cases. The idea is that the resulting equation and wavefunctions should give a physical description equivalent to the initial one. It is then claimed that this constraint determines what the value of $\xi$ is. However, nowhere in the derivation, the bounds of the integrals which define the actions are specified and the issues raised in the two previous paragraphs are ubiquitous. Moreover, as explained in our work, to conclude that a priori different Hilbert spaces of wavefunctions are equivalent, their inner products must be equal. However, inner products are not discussed in Ref. [10].